Problem 3-7

Freight car loadings over an 18-week period at a busy port are as follows:

Week Number Week Number Week Number
1 370 7 415 13 450
2 380 8 425 14 455
3 390 9 435 15 475
4 380 10 425 16 485
5 390 11 435 17 495
6 395 12 445 18 505


a.
Determine a linear trend line for expected freight car loadings. (Round your intermediate calculations and final answers to 2 decimal places.)

YˆY^ =  +  t

b. Use the above trend equation to predict expected loadings for Weeks 20 & 21. (Round your final answers to 2 decimal places.)

The forecasted demand for Week 20 is              and for Week 21 is              .

c. The manager intends to install new equipment when the volume exceeds 950 loadings per week. Assuming the current trend continues, in which week (at the earliest) should the loading volume reach that level? (Use the rounded answers, as required, from any previous part of this problem. Do not round any other intermediate calculations. Round your final answer to 2 decimal places.)

It should reach 950 loadings in Week              .
Answer #1

Let X denotes the Week and Y denotes the Number, Now compute the following Number Week X XYx²y x2 2 3 7 10 11 12 13 14 15 16

Let the linear regression curve Y = B. + Bt The least squares estimate of the regression coefficients B, and B, are computedY = 357.62 +7.68 Y_1 = 357.62 +7.68(21) = 357.62 +7.68(21) = 357.62 +161.28 = 518.9 Hence, the expected loading for weeks 20